A simple pendulum has a period of 2.5 s. What is its period if its length is made four times larger?
1
Expert's answer
2021-02-10T10:11:21-0500

The period of the pendulum can be written as follows:


"T=2\pi \sqrt{\dfrac{L}{g}}."

If the length of the pendulum is four times larger, then the period will be:


"T_{new}=2\pi \sqrt{\dfrac{4L}{g}}=2\cdot2\pi \sqrt{\dfrac{L}{g}}=2T,""T_{new}=2\cdot2.5\ s=5\ s."